OpenStudy (anonymous):

I've been going at this for hours now and simply can't grasp the concept. If you need a question answered, feel free to send them my way. Thanks! Use the substitution method to solve the following system of equations. x + 2y - z = 7 4x - y + 3z = -2 2x + 2y - z = 9 ANSWER CHOICES: (2, 1, 3) (2, 1, -3) (-2, 1, -3) (-2, -1, 3)

OpenStudy (shubhamsrg):

subtract 3rd and 1st eqn,, get value of x directly then its easier..

OpenStudy (anonymous):

I still do not understand

OpenStudy (shubhamsrg):

what do you get when you subtract 3rd and 1 st eqn ?

OpenStudy (anonymous):

OH, -2

OpenStudy (shubhamsrg):

thats on RHS,, what will be your LHS ?

OpenStudy (anonymous):

So it's either C or D, but what else do I have to figure out?

OpenStudy (shubhamsrg):

nops..you did mistake..check again..

OpenStudy (anonymous):

Come again? Not familiar

OpenStudy (anonymous):

7 - 9 = -2

OpenStudy (shubhamsrg):

(2x + 2y - z)-(x + 2y - z) = 9 - 7 x = ?

OpenStudy (anonymous):

Not sure what x equals, that's what I need to find

OpenStudy (anonymous):

What do I substitute in place of x?

OpenStudy (shubhamsrg):

you dont have a good conceptual base / understanding do you ?

OpenStudy (anonymous):

I obviously don't which is why I'm seeking for assistance

OpenStudy (anonymous):

You don't have to be an a**hole though

OpenStudy (preetha):

InsanelyChaotic, that was uncalled for. He is only trying to figure out how to best help you

OpenStudy (anonymous):

I viewed it as an insult

OpenStudy (anonymous):

My apologies if otherwise